### v1

# PEPT: Expert Finding Meets Personalized Pre-training

Abstract.
Finding appropriate experts is essential in Community Question Answering (CQA) platforms as it enables the effective routing of questions to potential users who can provide relevant answers. The key is to personalized learning expert representations based on their historical answered questions, and accurately matching them with target questions. Recently, the application of Pre-trained Language Models (PLMs) have gained significant attraction due to their impressive capability to comprehend textual data, and are widespread used across various domains. There have been some preliminary works exploring the usability of PLMs in expert finding, such as pre-training expert or question representations. However, these models usually learn pure text representations of experts from histories, disregarding personalized and fine-grained expert modeling. For alleviating this, we present a personalized pre-training and fine-tuning paradigm, which could effectively learn expert interest and expertise simultaneously. Specifically, in our pre-training framework, we integrate historical answered questions of one expert with one target question, and regard it as a candidate aware expert-level input unit. Then, we fuse expert IDs into the pre-training for guiding the model to model personalized expert representations, which can help capture the unique characteristics and expertise of each individual expert. Additionally, in our pre-training task, we design: 1) a question-level masked language model task to learn the relatedness between histories, enabling the modeling of question-level expert interest; 2) a vote-oriented task to capture question-level expert expertise by predicting the vote score the expert would receive. Through our pre-training framework and tasks, our approach could holistically learn expert representations including interests and expertise. Our method has been extensively evaluated on six real-world CQA datasets, and the experimental results consistently demonstrate the superiority of our approach over competitive baseline methods.  [^1] This work is an extended and revised version of research that appeared at EMNLP 2022 (Peng and Liu, 2022).
Expert Finding, Personalization, Community Question Answering, Recommender Systems, Pre-trained Language Model
copyright: nonecopyright: acmcopyrightjournalyear: 2023doi: 10.1145/1122445.1122456journal: TOISjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Information systems Recommender systems
1. INTRODUCTION
Recently, the popularity of online CQA websites based on user-generated content has been continuously increasing (Zhu et al., 2011b; Srba and Bielikova, 2016). For instance, some large websites such as Zhihu, Stack Exchange, Quora provide convenient bridges to exchange knowledge and have successfully garnered a substantial user base (Hu et al., 2019; Yuan et al., 2020b; Hu et al., 2021; Li et al., 2022). These are websites where users can pose questions or offer responses related to topics of their interest or expertise (Fang et al., 2016; Peng et al., 2022a). Due to the large participation, some questions remain unanswered for extended periods, resulting in unsatisfactory response times for askers (Zhao et al., 2017; Yuan et al., 2020b). Therefore, routing questions to suitable experts who can provide satisfactory answers (Chang and Pal, 2013; Zhao et al., 2014) is crucial to CQA websites. Expert finding on the CQA websites can effectively recommend questions to suitable experts (Zhu et al., 2014) capable of providing high-quality answers. Then, the askers could receive answers quickly (Liu et al., 2022a). This approach has recently garnered considerable attention in both academic and industry.
A key challenge in expert finding is the accurate matching of candidate experts with target questions (Li et al., 2019; Zhang et al., 2020). Thus, it is crucial to learn superiority expert representations to achieve precise question-expert matching. Existing expert finding methods mostly learn expert representations from limited supervised data (e.g., experts’ previously answered questions) and subsequently compute the expert-question matching scores. For example, Pal et al. (Pal et al., 2011) model users from their early historical activities, then use classification and ranking models to find potential experts. NeRank (Li et al., 2019) considers the question raiser’s profile and question content to learn entities representations with a network embedding method, which could help capture personalized expert representations. Peng et al. (Peng et al., 2022a) design a multi-view learning paradigm (PMEF) based on the question title, body and tag information. It could learn multi-view question/expert representations separately, and then compute question-expert relatedness for identifying experts.
In recent years, the language modeling capability of PLMs (e.g., BERT (Devlin et al., 2018)), has demonstrated remarkable improvements. Pre-training the Transformer Encoder (Vaswani et al., 2017), on large unsupervised corpora enables models to learn more generalized knowledge representations. Extensive research has shown that pre-trained models are capable of capturing intricate sequence patterns from large amounts of raw data. Then it can transfer this knowledge to a range of downstream NLP tasks, especially in situations where limited training data is available (Brown et al., 2020; Raffel et al., 2020). Inspired by this, researchers been exploring the feasibility of pre-training technology for finding suitable experts (Liu et al., 2022b; Peng and Liu, 2022). For example, one preliminary work ExpertPLM (Peng and Liu, 2022) designs a tailored pre-training model for modeling expert based on the historical answered questions, which could help learn expert representations from unsupervised data.
Although previous studies have demonstrated effectiveness, we argue that they still have several limitations that need to be addressed:
1) Interest Modeling. Experts’ interests reflect their willingness to answer questions, which could be captured from their historical answered questions. Existing Pre-trained Language Model (PLM)-based expert finding methods typically concatenate all questions an expert has answered and employ the Masked Language Modeling (MLM) task for modeling. However, this approach only captures expert interest based on word-level information and neglects question-level interest. In fact, expert interest is better represented by question-level semantics rather than individual keywords. Therefore, how to capture the question-level interest is important during pre-training.
2) Expertise Modeling. In fact, the ability of an expert to answer a target question plays a crucial role in identifying a suitable expert. For instance, one expert has answered questions “Was public transportation free in the Soviet Union?” and “Is there any record of ancient Israelite contact with Indo-Europeans?”, which could reflect his interest in “Soviet Union” and “Indo-Europeans”. However, the vote scores received for two different questions can vary significantly (e.g., $+26$ and $-5$), indicating variations in the expert’s expertise across different questions. Nevertheless, most existing pre-trained language model-based methods struggle to capture an expert’s ability to answer a specific question.
3) Personalized Modeling. It is intuitive that different experts exhibit diverse personalized characteristics. For example, there is one question “Why did Hitler … Soviet Union … fighting the United Kingdom?” waiting for answers. If one Expert is more concerned about “Hitler” than “Britain”, and another expert is more concerned about “Britain” than “Hitler”, their concerns on the question are different. Hence, the question differs in usefulness for them when learning their representations. However, existing PLM-based methods usually ignore these essential personalized characteristics.
Hence, for alleviating the above limitations, we propose to enhance the expert finding via a Personalized Expert-level Pre-Training language model (named PEPT), which could personalized pre-train expert representations before the downstream task expert finding. Our approach addresses the aforementioned limitations as follows:
1) Interest modeling. To enhance the pre-trained model’s understanding of experts’ interests, we integrate the expert’s historical answered questions as a source of expert-level interest information and utilize it as input for the model. Furthermore, we design a question-level pre-training strategy to capture underlying and inherent relationships between different historical questions. Compared to the traditional masked language model task for pre-training expert interests, our approach could learn more fine-grained expert interest features.
2) Expertise modeling. One way to assess an expert’s ability to answer questions is by using the vote scores, which indicate the level of CQA user appreciation for a particular answer. In other words, higher vote scores generally indicate higher levels of expertise associated with the user providing that answer. Thereby, to learn an expert’s fine-grained abilities of answering questions, we incorporate his/her historical answered question input embedding with corresponding encoded vote scores. Additionally, given one target question, we introduce a vote-oriented task to predict the vote score an expert would receive, empowering the model a better perception of an expert’s ability to answer the target question.
3) Personalized modeling. To further enhance personalized expert characteristics learning, we introduce the unique Expert ID into the model, which could indicate different expert personalized characteristics. One simplistic approach is to translate numerical expert IDs into their corresponding word embeddings, but this method can lead to excessive vocabulary size. Inspired by continuous/soft prompts (Lester et al., 2021), we directly input Expert ID vectors into the model by incorporating them with the expert-level information, which enhances the model to achieve personalized pre-training.
In a word, we propose to unify heterogeneous information into the pre-training framework to provide rich foundations for learning expert representation. Furthermore, we design a multi-task pre-training paradigm for guiding the model to capture fine-grained expert interest and expertise.
During finding experts, we construct the model input like that in pre-training for ensure the input consistency, then compute the expert-question matching score. This approach enables us to directly identify suitable experts through a fine-tuning way.
In a word, our proposed model makes the following main contributions:
  • We design a personalized pre-training language model paradigm in expert finding, named PEPT, aiming to pre-train personalized and fine-grained expert representations, and then conduct the expert finding in a fine-tuning way. To our knowledge, this is the first study to design a personalized pre-training paradigm for expert finding, enabling to capture more comprehensive expert features derived from unsupervised data.
  • We unify the previous questions, corresponding vote scores, and personalized information in the pre-training stage to model experts. Beneficial from this framework, we can directly conduct the expert finding based on the pre-trained model without gaps. Furthermore, during pre-training, we design a question-level MLM task and a vote-oriented task, which allows to better model the fine-grained expert characteristics.
  • Through experiments on six real-world community question answering website datasets, we can find PEPT outperforms the recent baselines. These results demonstrate the superiority of PEPT in enhancing expert finding. Additionally, our further analysis showcases the effectiveness of different model modules and the potential of pre-trained representations for improving expert finding.
Differences between this work and the conference version. This work is an extended and revised version of research that appeared at EMNLP 2022 (Peng and Liu, 2022). The current version differs significantly from its preliminary conference version in the pre-training paradigm. Specifically,
  1. (1)
    We introduce the personalized information of experts into the expert pre-training, which enables the model to learn the unique representations of experts effectively. However, the previous version did not take into account such personalization;
  2. (2)
    This work designs a more fine-grained expert pre-training architecture compared with our previous version. Specifically, our conference version only considers modeling the coarse-grained expert-level characteristics, and our extended version designs two question-level pre-training tasks, which could empower the model ability to capture more fine-grained interests and capabilities;
  3. (3)
    We integrate target questions in the expert-level information, which can help the model pre-train candidate-aware expert representations. Another advantage of this design is that it can eliminate the gap between the pre-training and expert finding, i.e., there is no need to design additional encoders during fine-tuning like the previous version but can be unified in one model;
  4. (4)
    In addition, this paper conduct various of experiments aimed at corroborating the effectiveness of our pre-training paradigm.
2. RELATED WORK
We provide a concise literature review on three key topics relevant to our proposed method: Pre-training for Natural Language Processing (NLP), Pre-training for Recommender Systems (RS), and Expert Finding in this section.
Table 1. Characteristics of Different Expert Finding Methods.
Method Characteristic Interest Expertise Personalized Pre-training
Doc2Vec $\times$ $\times$ $\times$ $\times$
CNTN (Qiu and Huang, 2015) $\times$ $\times$ $\times$ $\times$
NeRank (Li et al., 2019) $\surd$ $\times$ $\surd$ $\times$
TCQR (Zhang et al., 2020) $\surd$ $\times$ $\surd$ $\times$
RMRN (Fu et al., 2020) $\surd$ $\times$ $\times$ $\times$
UserEmb (Ghasemi et al., 2021) $\surd$ $\times$ $\surd$ $\times$
PMEF (Peng et al., 2022a) $\surd$ $\times$ $\surd$ $\times$
EFHM (Peng et al., 2022b) $\surd$ $\times$ $\surd$ $\times$
ExpertBert (Liu et al., 2022b) $\surd$ $\times$ $\times$ $\surd$
PEPT $\surd$ $\surd$ $\surd$ $\surd$
2.1. Pre-training for Natural Language Processing
Recently, various of Pre-Training Models (PTMs) have demonstrated the efficacy of pre-training on large corpora and the capability to acquire general language knowledge, which can be effectively transferred to various downstream NLP tasks. (Qiu et al., 2020). Earlier word embedding methods, such as Word2Vec (Mikolov et al., 2013), Glove (Pennington et al., 2014) learned good word embeddings by capturing word co-occurrence information. These embeddings derived from these models could significantly improve various NLP missions (e.g., text classification (Lilleberg et al., 2015), named entity recognition (Lample et al., 2016)) by capturing the semantic meanings of words. However, these models are context-independent and predominantly trained using shallow models, which may restrict their effectiveness in specific scenarios that necessitate a more profound comprehension of language context.
To overcome these limitations, various of pre-training methods (Brown et al., 2020) based on the Transformer architecture (Vaswani et al., 2017) have been developed, showing their significant improvements in learning contextual language representations. These approaches, such as BERT (Devlin et al., 2018), MASS (Song et al., 2019), and BART (Lewis et al., 2020), mainly pre-train model parameters based on various neural network (e.g., transformer encoder, transformer decoder) and language modeling objectives. Then fine-tune the model parameters with downstream supervised data. By employing pre-training and fine-tuning paradigms, it can proficiently utilize extensive corpora to acquire comprehensive language knowledge. At the same time, the model could integrate task-specific information during the fine-tuning phase. Thus, these models have the potential to enhance the performance of diverse downstream tasks.
However, most existing PLMs are carefully designed at the corpus-level to capture universal language knowledge, limiting their ability to model experts. Therefore, it is nontrivial to directly employ the original pre-training models and the CQA corpus (e.g., question titles) for learning expert representations.
2.2. Pre-training for Recommender Systems
In recent times, pre-training technology has been increasingly used in recommendation models to enhance the performance of recommendation systems. Specifically, these methods (Kang and McAuley, 2018; Sun et al., 2019) leverage pre-training techniques to learn item co-occurrence information, which can be applied to improve the recommendation. For example, Cheng et al. (Cheng et al., 2021) proposed CLUE. It employed contrastive learning as its optimization objective during pre-training, with a focus on optimizing sequence-level representations. UniSRec (Hou et al., 2022) captured more transferable representations from the item texts for modeling users and items, which could effectively generalize to new recommendation scenarios. Then it fine-tuned the pre-trained model for conducting recommendation and user-oriented downstream tasks, which obtained better performance. Furthermore, many researchers have aimed to capture universal user embeddings that could be applied to various user-oriented downstream tasks (Yuan et al., 2020a; Xiao et al., 2021). For example, Xiao et al. (Xiao et al., 2021) proposed UPRec, which integrated heterogeneous user information and employed two pre-training tasks to learn user-aware representations.
However, unlike the above sequential recommendation tasks, expert finding on CQA websites is always faced with the challenge of cold-start (i.e., the target question is always a new question). Moreover, it has unique characteristics that require specialized approaches, such as expertise modeling of experts, which aims to reflect the ability of experts to provide accurate and helpful answers. As a result, developing effective pre-training approaches for expert finding requires careful consideration of these unique characteristics.
2.3. Expert Finding
The primary objective of expert finding in the context of CQA websites is to locate individuals who possess the necessary expertise and knowledge to provide satisfactory responses to user questions (Zhu et al., 2011a; Yuan et al., 2020b). The existing body of research in this domain can be broadly classified into two major categories: traditional expert finding methods and expert finding methods based on deep learning techniques.
Traditional methods often relied on feature engineering (Robertson and Zaragoza, 2009; Cao et al., 2012; Pal et al., 2012) or topic modeling techniques (Riahi et al., 2012; Ji and Wang, 2013; Yang et al., 2013) to learn question and expert representations. Afterwards, the model utilized the question and expert representations for computing the relevance scores and routing questions to suitable experts. For instance, Zhou et al. (Zhou et al., 2012) captured local-/global-level characteristics based on questions, users, and relations between them, then these features were fed to the SVM (Smola and Schölkopf, 2004) for finding suitable experts. Chang et al. (Chang and Pal, 2013) employed LDA (Blei et al., 2003) to model the topic information from the question contents and took that information as the question representations for recommending experts. Neshati et al. (Neshati et al., 2017) introduced two learning algorithms that considered the temporal-aware aspects of expert finding. They combine factors such as topic similarity to predict a user’s likelihood of becoming an expert. Deep learning-based methods (Huang et al., 2013; Qiu and Huang, 2015; Wang and Nyberg, 2015) leveraged neural networks to model questions and experts, then computed the question-expert relevance for finding experts (Li et al., 2019; Fu et al., 2020; Ghasemi et al., 2021). For example, TCQR (Zhang et al., 2020) employed the pre-trained BERT to learn question semantic information and captured the answerer’s features based on the question semantic and temporal information for expert finding. Qian et al. (Qian et al., 2022) designed multi-perspective metapath-based network and considered vote scores for question routing, but they omitted the temporal interest of the expert. HQExpert employed (Liu et al., 2022c) different granularity semantic information and temporal interest to accurately identify experts and improve accuracy, but they could not consider personalized characteristics of the expert and the varying expert-question relationships.
In recent years, PLMs have emerged as a powerful tool for improving the performance of various NLP tasks, including textual similarity, etc. (Zhang et al., 2021; Xu et al., 2023). This is primarily due to their superior ability to understand and analyze complex textual data, enabling them to capture more nuanced features and relationships within language. Some works have focused on leveraging the capabilities of PLMs to develop advanced expert finding techniques (Liu et al., 2022b; Peng and Liu, 2022). For instance, Liu et al. (Liu et al., 2022b) designed a semi-supervised pre-training paradigm for expert finding, which could incorporate downstream tasks into the pre-training and improve the model’s ability to utilize limited data.
Based on the existing methods discussed above, we have identified several key characteristics of advanced expert finding techniques. These include interests, expertise, personalized, and pre-training. We list them in Table 1. Our proposed method incorporates all the aforementioned key characteristics of expert finding, which could pre-train more comprehensive personalized characteristics from unsupervised data for improving expert finding.
3. PROBLEM DEFINITION
We first formal definition of the expert finding, then give some important notations and explanations used throughout the remainder of this paper.
In the expert finding task, the main goal is to match a target question, denoted as $q^{t}$, with the most appropriate expert who is capable of providing a relevant and accurate answer. Assuming that there is a candidate expert set $C^{u}=\{c^{u}_{1},\cdots,c^{u}_{M}\}$ and a target question $q^{t}$ to be answered respectively, where $M$ is the number of experts. In the set $C^{u}$, the candidate expert $c_{i}^{u}$ is associated with historically answered questions, denoted as $Q_{i}^{u}$. The $Q_{i}^{u}$ includes $\{q_{1},\cdots,q_{n}\}$, where $n$ is historical question number. Each question is represented as a question title including a sequence of words. The vote scores obtained for each previously answered question of the expert can be presented as $V^{u}_{i}=\{v_{1},\cdots,v_{n}\}$. It is important to note that, the individual who provides the “accepted answer” for a given question is considered the expert or the ground truth expert. Each question has only one “accepted answer”.
The main notations are summarized in Table 2.
Table 2. Notations and Explanations.
Notations Definitions
$C^{u}$ Candidate expert set
$q^{t}$ Target question
$c^{u}_{i}$ $i$-th candidate expert
$M$ Number of experts
$Q^{u}_{i}$ $i$-th expert historical answered questions
$V^{u}_{i}$ $i$-th expert historical obtained vote scores
$n$ Number of expert historical questions
$q_{n}$ $n$-th historical question of the expert
$v_{n}$ $n$-th historical vote score of the expert
$d$ Representation vector dimension
$h$ The number of transformer heads
$t_{num}$ The token number of one input line
$\mathcal{V}$ The vocabulary in pre-training
$\mathcal{E}$ The model input information
$tk$ The input token
$\mathbf{R}_{t}$ The input representation matrix
$\mathbf{R}_{token}$ Token embedding matrix
$\mathbf{R}_{seg}$ Segment embedding matrix
$\mathbf{R}_{pos}$ Position embedding matrix
$\mathbf{R}_{v}$ The expert vote score embedding matrix
$\mathbf{R}_{in}$ The final expert-level embedding matrix
$\mathbf{C}^{u}$ Expert embedding matrix
$\mathbf{c}^{u}_{i}$ $i$-th expert embedding
$\mathbf{e}$ The expert pre-trained representation
$\mathbf{Q}$ Attention query parameter
$\mathbf{K}$ Attention key parameter
$\mathbf{V}$ Attention value parameter
$E_{c}$ The matching score predicted by the model
$\hat{E}_{c}$ Expert-question matching ground truth
$S_{v}$ Vote score predicted by the model
$\hat{S}_{v}$ Vote score ground truth
$\mathcal{L}_{ql}$ Question-level interest loss
$\mathcal{L}_{vs}$ Vote-oriented loss
$\mathcal{L}_{wm}$ Masked language model loss
4. PROPOSED METHOD
In this part, we provide an introduction to our method PEPT. Our method including two stages: pre-training expert representations and fine-tuning with the expert finding task. These stages are illustrated in the Fig 1 and Fig 2.
In the pre-training, we incorporate questions answered by one expert historically and the corresponding target question for constructing target-aware expert-level input units. Besides, we introduce the expert’s personalized information to guide personalized expert pre-training, which could help capture the expert’s unique characteristics. For pre-training, we design the question-level MLM and the vote-oriented tasks for leading the question-level expert interest and expertise learning. The question-level MLM task is employed to forecast the masked question by taking into account the contextual questions, which could better understand the historical question relatedness for modeling interests. The vote-oriented task integrates vote scores from the corresponding historical questions (excluding the target question) and enables the model to predict the expert’s likely vote score for one target question.
Beneficial from our designed pre-training paradigm, the inputs to the model are very similar between pre-training stage and downstream fine-tuning stage. And hence, we can directly employ the pre-trained model for conducting expert finding task without any model architecture modification. The subsequent subsections describe each phase in detail.
Figure 1. PEPT Personalized Pre-training Framework. We unify heterogeneous information into the pre-training model to provide rich fundamental information for learning expert representations. During pre-training, we additionally design a question-level masked language model and a vote-oriented task to empower the model abilities to capture expert interest and expertise.
4.1. Pre-training Stage
The main difference between this work and the conference version lies in the design of the pre-training architecture. In this work, we introduce personalized expert information into the model through the soft prompt to achieve personalized pre-training. Via adopting this approach, the model becomes capable of capturing and understanding the distinct attributes of each individual expert. Additionally, by designing specific question-level pre-training tasks, the model gains the ability to capture more fine-grained expert interests and capabilities, when facing with different target questions.
In the following sections, we present the PEPT from three perspectives: (1) Input layer for introducing the input information constructed for the model; (2) Model architecture for introducing the model components; (3) Pre-training task for introducing our specially tailored training tasks for expert finding.
4.1.1. Input Layer
Different from traditional language knowledge pre-training in Nature Language Progressing, which concentrates on corpus-level information input, we aim to empower the BERT abilities to personalize capture expert representations. We introduce the input layer from the three separate aspects (i.e., interest, expertise, and personalized).
Interest Information
In CQA websites, the user’s interests can be reflected from the questions he/she has answered historically. And hence, as shown in Fig 1, we construct a unified expert-level sequence by directly concatenating the words from the expert’s previously answered questions, allowing the model to capture the expert’s interest information. Besides, experts’ representations should be different when faced with different target questions. Therefore, we further incorporate the target question information with the expert-level interest information.
For prompting our model to clearly identify the target question and each previously answered question during pre-training, we design two special tokens and add them to the input sequence.
  • [HSEP]. This special token is designed to separate different historical questions from one expert.
  • [SEP] (Devlin et al., 2018). We utilize that to distinguish target questions and expert’s historical questions.
And hence, the input information constructed for the $i$-th expert can be formulated as:
(1) $\begin{split}\mathcal{E}=[\underbrace{[{tk}_{1}],[{tk}_{2}],[{tk}_{3}]}_{q^{t}},\mathtt{[SEP]},\\ \underbrace{\underbrace{[{tk}_{1}],[{tk}_{2}]}_{q_{1}},\mathtt{[HSEP]},\underbrace{[tk_{1}],[tk_{2}],[tk_{3}]}_{q_{2}}}_{c^{u}_{i}},\mathtt{[SEP]}]\end{split}$
Subsequently, taking into account the extensive linguistic knowledge already captured by the original pre-trained BERT weight (Devlin et al., 2018), we employ that to initialize the model parameters. Noted that the segment ID is always to 00, since we don’t employ the Next Sentence Prediction (NSP) task. And the input representation matrix $\mathbf{R}_{t}\in\mathcal{R}^{t_{num}\times d}$ ($t_{num}$ is the number of tokens) is generated by summing its corresponding embeddings:
(2) $\mathbf{R}_{t}=\mathbf{R}_{token}+\mathbf{R}_{seg}+\mathbf{R}_{pos}\ ,$
where $\mathbf{R}_{token}$ is the token embedding, and $\mathbf{R}_{seg}$ is the segment embedding, and $\mathbf{R}_{pos}$ is the position embedding.
Expertise Information
In general, the vote scores received by an expert can indicate their expertise in answering different questions. A higher vote score reflects a stronger professional expertise in addressing such questions. To effectively measure expert abilities in different question fields, we incorporate vote scores into our pre-training process.
Specifically, we pre-process vote scores (introduced in Implementation Details) of questions and employ an embedding matrix to map each vote score as an embedding vector (with $d$ as dimension). To accurately capture the expert’s expertise in a specific question, we ensure that the dimension of each vote score embedding matches the dimensionality of the corresponding historical question. In other words, the length of one question token is $x$, and the corresponding vote score embedding matrix is $x\times d$. Then we stack these vote score embeddings as a matrix $\mathbf{R}_{v}\in\mathcal{R}^{t_{num}\times d}$. Afterwards, we integrate $\mathbf{R}_{v}$ with the input representation $\mathbf{R}_{t}$ as follows:
(3) $\mathbf{R}_{in}=\mathbf{R}_{t}+\mathbf{R}_{v}\ .$
It is noted that we employ the special token [PAD] to replace the corresponding vote score of the target question for prediction.
Personalized Information
Different from NLP, expert finding further emphasize personalization. User ID embeddings are commonly used in recommendation models and could be considered as the identifying information of individual users (Rendle et al., 2009). Hence, we introduce the expert ID into the pre-training, which could reflect unique expert characteristics and help personalized pre-training. A simple and intuitive approach is to consider expert IDs as distinct word tokens and incorporate them into the pre-trained model’s vocabulary. However, considering potentially millions of experts on the CQA websites, this would lead to an excessively large word vocabulary, making pre-training time-consuming. It’s worth noting that the expert ID doesn’t need to be in word form and can be represented by a vector embedding created through random initialization or other techniques. This type of processing method can be formally termed soft prompt. Therefore, we treat expert IDs as additional prompting tokens rather than adding them to the vocabulary.
Specifically, we prepare one token embedding matrix $\mathbf{C}^{u}\in\mathcal{R}^{M\times d}$, where $M$ is the number of experts. Then, $i$-th expert’s vector representation is retrieved by:
(4) $\mathbf{c}^{u}_{i}=\mathbf{C}^{u}f(i),$
where $\mathbf{c}^{u}_{i}\in\mathcal{R}^{1\times d}$, and $f(i)\in{\{1,0\}}^{M}$ represents a one-hot vector, where the $1$ indicates the position of the expert’s vector in $\mathbf{C}^{u}$. Then, we concatenate the ID embedding of the expert at the begin of the corresponding $\mathbf{R}_{in}$, which is denoted as $[\mathbf{c}^{u}_{i};\mathbf{R}_{in}]$ with $\mathcal{R}^{(t_{num}+1)\times d}$ as dimension. This approach provides the foundation for the model to personalized pre-train expert interest and expertise across various questions.
4.1.2. Model Architecture
BERT (Devlin et al., 2018) contains multiple bidirectional Transformer encoders. Within each layer, self-attention and feed-forward operations are conducted to capture contextual information and generate contextualized representations of the input tokens.
Self-Attention Layer
Self-attention plays a vital role as a fundamental building block within the Transformer architecture. By enabling each token to attend to all other tokens in the sequence, BERT effectively captures contextual relationships between the tokens. The self-attention function is described as follows:
(5) $Att(\mathbf{Q},\mathbf{K},\mathbf{V})=Softmax(\mathbf{Q}\mathbf{K}^{T}/\sqrt{d})\mathbf{V}\ .$
Besides, the BERT employs multi-head attention, where self-attention is performed in parallel multiple times. This function MH($\cdot$) allows the model to capture different types of contextual relationships, which can be denoted as:
(6) $MH(\mathbf{R}_{in})=[{head}_{1};\cdots;{head}_{h}]\mathbf{W}\ ,$
(7) ${head}_{i}=Att(\mathbf{R}_{in}\mathbf{W}^{q}_{i},\mathbf{R}_{in}\mathbf{W}^{k}_{i},\mathbf{R}_{in}\mathbf{W}^{v}_{i})\ ,$
where $\mathbf{W}^{q}_{i},\mathbf{W}^{k}_{i},\mathbf{W}^{v}_{i}\in\mathcal{R}^{d\times\frac{d}{h}}$ and $\mathbf{W}\in\mathcal{R}^{d\times d}$ are learnable parameters. Through this function, the $\mathbf{R}_{in}$ is transformed to $\mathbf{H}\in\mathcal{R}^{t_{num}\times d}$ ($t_{num}$ is the token number), which capture the contextual relationships.
Feed-Forward Layer
Then, we utilize feed-forward layers within each transformer encoder layer to apply non-linear transformations to the representations captured by the self-attention mechanism. Given the input matrix $\mathbf{H}$, the calculation of this function is defined as:
(8) $FFN(\mathbf{H})=RELU(\mathbf{H}\mathbf{W}^{f}_{1}+b^{f}_{1})\mathbf{W}^{f}_{2}+b^{f}_{2}\ ,$
where $\mathbf{W}^{f}_{1}$, $\mathbf{W}^{f}_{2}$ and $b^{f}_{1}$, $b^{f}_{2}$ are learnable parameters. More details please refer to (Vaswani et al., 2017).
Figure 2. PEPT Personalized Fine-tuning. It is noted that [+39] and [-4] represent different question vote scores. The pre-trained encoder is the pre-trained expert weight from PEPT Personalized Pre-training.
4.1.3. Pre-training Task
We introduce our designed pre-training tasks in this part, which are the core mechanisms that encourage PLMs to model experts’ interest and expertise across various questions.
Interest Pre-training
In fact, the interest of experts could be reflected in the model input sequence that we design. As illustrated in Section 1, modeling the question-level semantic information is important for expert interest modeling. However, existing preliminary works could only capture interests from word-level information rather than question-level semantic information. As expert histories are diverse and harder to predict, different from the MLM task, we propose an additional pre-training task called question-level masked language model to capture question-level interests.
Specifically, given the example illustrated in Fig 1, inspired by SpanBERT (Joshi et al., 2020), we mask all title words of the question $q_{2}$, then predict the token [Jews] in the $q_{2}$ based on the contextual semantic information. And the loss function is:
(9) $\mathcal{L}_{ql}=-logp(\texttt{Jews}|x_{11},x_{16},p_{14}),\texttt{Jews}\in[1,2,\cdots,\lvert\mathcal{V}\rvert]\ ,$
where $x_{11}$ and $x_{16}$ represent the contextual historical question semantic information, and $p_{14}$ represents the position of the token that is predicted by the model.
Expertise Pre-training
As denoted above, the vote score the expert received could reflect the expertise in answering one question. Generally, higher vote scores indicate higher expertise, and the answer would be more satisfied the CQA users. Therefore, we design a vote-oriented task to pre-train the model and enhance its capacity to discern the expert’s expertise in addressing the target question. Specifically, given the instance shown in Fig 1, the model output embedding from the expert ID can be viewed as an expert-level representation $\mathbf{e}\in\mathcal{R}^{1\times d}$. Hence, given a target question, we directly employ $\mathbf{e}$ to predict the vote score that the expert would receive, which is defined as:
(10) $S_{v}=\mathbf{W}_{v}(\mathbf{e^{T}})+\mu_{v}\ ,$
where $\mathbf{W}_{v}\in\mathcal{R}^{1\times d}$ is the model parameter and $\mu_{v}$ denotes the bias. As the vote prediction is a regression problem, we employ the square loss function to predict the received vote score for pre-training our model, which is formulated as follows:
(11) $\mathcal{L}_{vs}=\sum_{\mathbf{\Omega}}(\hat{S}_{v}-S_{v})^{2}\ ,$
where $\mathbf{\Omega}$ denotes the all instance in the pre-training stage, and $\hat{S}_{v}$ is the ground truth vote score.
Additionally, acknowledging the potent capacity of the original MLM task (Devlin et al., 2018) to train PLMs, we incorporate this task to acquire linguistic knowledge in CQA websites. The MLM task randomly masks certain words and subsequently leverages bidirectional contextual information to reconstruct the input sequence. The formal definition is:
(12) $\mathcal{L}_{wm}=-\sum_{i=1}^{\mathcal{N}}logp(n=n_{i}|\theta),n_{i}\in[1,2,\cdots,\lvert\mathcal{V}\rvert],$
where $\mathcal{N}$ is the masked token number, and $\mathcal{V}$ is the vocabulary.
4.1.4. Model Learning
During pre-training, the PEPT model is collectively trained in three tasks, and the loss function used is:
(13) $\mathcal{L}_{pt}=\mathcal{L}_{ql}+\mathcal{L}_{vs}+\mathcal{L}_{wm}\ .$
Through the integration of these loss functions, our PEPT model effectively pre-trains personalized expert-level representations that encompass not only the interests but also the expertise of the experts.
4.2. Fine-tuning Stage
In this section, we perform the expert finding task using a fine-tuning approach. As depicted in Fig 2, we combine the previously answered questions by the expert with the target question and input the resulting sequence directly into the pre-trained model, which is denoted as $[\mathbf{c}^{u}_{i};\mathbf{R}_{in}]$. By adopting this approach, we ensure that the inputs during the fine-tuning stage remain consistent with those used during the pre-training phase. This consistency helps minimize the disparity between the pre-training and fine-tuning stages.
Next, we add a linear layer to the special embedding $\mathbf{e}$ (derived from the ID embedding) to predict the expert-question matching probability $E_{c}$, denoted as:
(14) $E_{c}=\mathbf{W}_{c}(\mathbf{e}^{T})+\mu_{c}\ ,$
where $\mathbf{W}_{c}$ and $\mu_{c}$ are model parameters.
During fine-tuning, we follow previous works (Li et al., 2019; Peng et al., 2022a) and utilize the negative sampling technology (Huang et al., 2013) to generate $k$ negative samples. Then, we employ the cross-entropy loss for optimizing, which can be expressed as follows:
(15) $E_{c}=\frac{\exp(E_{c})}{\sum_{j=1}^{K+1}\exp(E_{j})}\ ,Loss=-\sum_{c=1}^{k+1}\hat{E}_{c}log(E_{c})\ ,$
where $\hat{E}_{c}$ is the ground truth and $E_{c}$ is the normalized probability predicted by methods.
5. EXPERIMENTAL SETTINGS
Table 3. Statistical Details of Each Dataset.
Datasets # density (%) # questions # answerers # answers # avg.title length
Es 0.0429 36,271 3,260 50,801 10.05
Gis 0.0435 50,718 3,168 70,034 9.75
Biology 0.2081 8,704 630 11,411 9.84
English 0.0468 46,692 4,781 104,453 9.68
Electronics 0.0585 56,614 3,084 102,214 9.28
CodeReview 0.0696 36,947 2,242 57,622 7.27
To evaluate the performance of PEPT, we conduct a series of comprehensive experiments on six CQA datasets. To begin, we provide an introduction to the datasets employed in our paper. Afterwards, we give the introduction to the implementation details of our experiment. Then, we introduce the baseline methods and evaluation metrics used in this paper.
5.1. Datasets
For our experiments, we use a dataset comprising 103,005 experts extracted from StackExchange [^2] https://archive.org/details/stackexchange. Each expert corresponds to his/her historical answered questions. This dataset is used for pre-training expert representations in PEPT. We evaluate the effectiveness of the pre-trained model in specialized domains by selecting six distinct domains: English, Biology, Es, Electronics, Gis, and CodeReview. Within each specific domain, we organize a collection of questions. Each question is accompanied by a sequence of title words that provide question semantic information. Additionally, each question has an “accepted answer” chosen from multiple answers provided by different experts. In the evaluation, we consider the expert who has provided the “accepted answer” as the ground truth. To construct candidate experts for questions, we follow the pre-processing method employed in previous research (Peng et al., 2022a). This set includes the original experts who previously answered the question, and randomly selected additional experts from the pool of answerers (one set contains 20 experts in total). Table 3 provides an overview of the statistical properties of the datasets, offering a summary of their key characteristics. To ensure a comprehensive evaluation, we split each dataset into three distinct subsets: training/validation/testing sets. The proportions of these subsets, in chronological order, are set to 80%, 10%, and 10% of the dataset, respectively. Noted that the validation set and the testing set do not appear in the pre-training data.
5.2. Implementation Details
Within this section, we present a comprehensive overview of the experimental settings and hyperparameters employed in our model. Next, we provide an introduction of the vote score pre-processing technique. It should be noted that all hyperparameters in our model were fine-tuned using the validation dataset.
Experimental Setting
Our personalized pre-training (PEPT) model, with approximately 110 million parameters, is based on the bert-base-uncased (Devlin et al., 2018). The batch size for both pre-training and fine-tuning is $16$. For alleviating over-fitting, we use dropout technology (Srivastava et al., 2014). Specifically, the dropout ratio is $0.2$ in pre-training and $0.1$ in fine-tuning. We optimize our model using the Adam optimization strategy (Kingma and Ba, 2015). We set the learning rate to $55e-$ 5 during pre-training and $71e-$ 7 during fine-tuning. The token embedding and expert representation dimensions are both $768$. The maximum training step during pre-training is 50,000. We conduct each experiment independently and repeated it three times, and report the average results. All experiments were carried out using the PyTorch framework, which served as the foundation for our implementation. We utilize a single GPU server with an Intel(R) Xeon(R) CPU@2.20GHz and two RTX-3090 GPUs, each with 24GB of memory.
Pre-processing of Vote scores
In this part, we introduce how to process the vote scores in our dataset. One of the key challenges we encountered is a large variance in the distribution of vote scores. For example, in the Biology dataset, we observe the maximum vote score is $+287$, while the minimum is $-8$, and the average vote score is $3.49$.
Therefore, to alleviate this situation, we employ a normalization approach. First, we translate the vote scores overall to eliminate negative values (i.e., all vote scores are added to the minimum value in the vote score set). Subsequently, we apply a logarithmic transformation (i.e., $ln(\cdot)$) to the vote scores to reduce the impact of excessive variance. This helps ensure that the scores are more evenly distributed across the dataset. To simplify data processing and ensure roughly equivalent numbers of scores within each score segment, we normalized the vote scores to integers ranging from $1$ to $5$. The normalization process is conducted as follows:
(16) $v_{min}=min(V^{u}),v_{max}=max(V^{u})\ ,$
where $v_{min}$ and $v_{max}$ represent the minimum score and the maximum score of the $V^{u}$ respectively.
(17) $v^{*}_{i}=round(\frac{4}{v_{max}-v_{min}}*(v_{i}-v_{min})+1)\ ,$
here, $v^{*}_{i}$ represents the normalized score of the $v_{i}$ in $V^{u}$, and $round$ denotes the rounding operator.
5.3. Comparison Methods
To ensure a comprehensive evaluation of the effectiveness of PEPT, we compare its performance against two different types of competitive baselines:
First, the neural expert finding methods that utilize various neural networks to learn expert and question representations, then use these representations for finding experts.
  • Doc2Vec utilizes the Doc2Vec technology to learn expert/question representation and identifies suitable experts based on their past experience in answering similar questions to the target question.
  • CNTN (Qiu and Huang, 2015) employs the CNN to model question semantic information and model experts, then calculate the question-expert relevance scores.
  • NeRank (Li et al., 2019) learns question, raiser and expert representations via a Heterogeneous Information Network (HIN) (Sun et al., 2011) algorithm, then uses the CNN network to find the expert for questions.
  • TCQR (Zhang et al., 2020) incorporates a temporal context-aware model that captures temporal patterns in experts’ answering behavior across multiple time scales.
  • RMRN (Fu et al., 2020) incorporated reasoning memory cells (RMCs) to capture the essence of the question and leveraged an attention mechanism to extract relevant information from the historical records of candidate experts
  • UserEmb (Ghasemi et al., 2021) utilizes a node2vec (Grover and Leskovec, 2016) for learning structural relationships among users and uses a word2vec (Mikolov et al., 2013) to model the semantic context of words, and integrates the social features and word features to improve the expert finding.
  • EFHM (Peng et al., 2022b) is a multi-grained hierarchical matching framework for expert finding in CQA websites, which could capture more fine-grained matching relationships and achieve superior performance.
  • PMEF (Peng et al., 2022a) designs a multi-view attentive expert finding method with a personalized mechanism, which could comprehensively model expert/question and precisely match them.
Second, the latest PLM-based methods model expert or question features using pre-trained language models, which outperforms above baselines.
  • ExpertBert (Liu et al., 2022b) is a preliminary pre-training framework for expert finding. The framework is designed to accurately model the question-expert matching, which could improve performance of expert finding.
  • ExpertPLM (Peng and Liu, 2022) designs an expert-specific pre-training paradigm for finding expert. The key feature of ExpertPLM is a reputation-augmented masked language model that is designed to model an experts’ interests and expertise based on their previously answered questions and reputations.
Some papers have already published their source code on Github, such as NeRank (Li et al., 2019), PMEF (Peng et al., 2022a), etc. For these models, we download the code and check the model details and parameters based on their original paper to ensure that the model is consistent with the original paper. For models without publicly available source code, such as RMRN (Fu et al., 2020), TCQR (Zhang et al., 2020), etc. We strictly follow technical details reported by the paper to reproduce the model and set model hyperparameters based on the hyperparameter settings reported in the paper.
5.4. Evaluation Metrics
For evaluating the model’s performance, we use the following three widely employed evaluation metrics:
  • Mean Reciprocal Rank (MRR): Measures models in identifying the most appropriate relevant expert from a candidate expert set. It is computed as the reciprocal of the rank assigned to the first relevant expert, which is denoted as:
    (18) ${\rm MRR}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{r_{i}}\ ,$
    where $N$ represents the total number of target questions and $r_{i}$ refers to the rank assigned to the most suitable expert within the candidate expert set for a specific question.
  • Precision@$K$: We measure the proportion of target questions where the true expert is ranked within the top-$K$ positions among the candidate answerers. In our experiments, we set $K$ as 1 and 3.
  • Normalized Discounted Cumulative Gain ($NDCG$@$K$(Järvelin and Kekäläinen, 2002): The metric takes into account both the relevance of the recommended experts and their respective positions within the candidate expert set. It is calculated by summing the discounted relevance scores of the recommended items and normalizing it by the ideal $DCG$ score. And the $NDCG$ is computed as: $NDCG$@$K$ $=\frac{{DCG}_{K}}{{IDCG}_{K}}$. The ${DCG}_{K}$ is computed as:
    (19) ${DCG}_{K}=\sum_{i=1}^{K}\frac{r_{i}}{log_{2}(i+1)}\ ,$
    where $r_{i}=1$ when the $i$-th expert is the ground truth. ${IDCG}_{K}$ represents the ideal discounted cumulative gain. And then $NDCG$@$K$ is computed as: $NDCG@K=\frac{{DCG}_{K}}{{IDCG}_{K}}$. We set $K=20$.
Overall, the above diverse metrics are used to assess the ranking accuracy of the expert recommended by an algorithm. A higher value for each of these metrics indicates better performance of the model in accurately predicting the most relevant expert.
6. EXPERIMENTAL RESULTS
Within this section, we showcase the outcomes of our comprehensive experiments and conduct a detailed analysis of PEPT. Specifically, (1) Performance Comparison compares the PEPT with 10 competitive baselines; (2) Ablation Study explores how different parts of the model affect model performance; (3) Pre-trained Expert Representation Analysis analyzes what the pre-trained expert representations benefit the expert finding. (4) Hyper-parameter Analysis analyzes the model performance with different model hyper-parameters. (5) Case Study understand PEPT’s advantages in finding experts more intuitively.
6.1. Performance Comparison
[…]
Dateset English Gis CodeReview
Method Metric MRR P@@@@1 P@@@@3 NDCG@@@@20 MRR P@@@@1 P@@@@3 NDCG@@@@20 MRR P@@@@1 P@@@@3 NDCG@@@@20
Doc2Vec 23.26 15.23 29.46 35.22 29.12 16.08 28.73 35.19 30.15 17.31 31.99 42.46
CNTN 29.68 18.37 39.36 42.25 40.12 23.12 40.15 42.01 35.53 20.13 50.77 50.35
NeRank 48.95 27.16 61.43 56.41 46.97 30.32 5577 58.36 49.47 30.89 60.55 61.54
TCQR 34.25 19.27 49.87 48.22 45.53 26.34 5489 56.37 42.53 21.12 53.82 58.39
RMRN 46.77 25.22 61.62 56.75 48.97 32.39 57.77 58.32 43.11 25.17 55.80 58.92
UserEmb 31.73 19.56 42.36 45.51 32.23 24.33 44.77 45.62 39.15 20.31 51.26 52.46
EFHM 50.87 31.66 62.16 57.65 47.28 32.31 57.68 59.05 49.22 30.25 60.58 59.77
PMEF 49.47 28.65 63.14 58.75 48.61 33.11 58.88 59.11 50.20 31.39 61.61 61.70
ExpertBert 52.01 32.32 64.97 62.38 51.70 35.37 60.66 62.65 51.23 32.87 62.65 62.48
ExpertPLM 52.50 31.68 65.12 61.17 50.85 34.56 60.58 61.78 51.15 32.63 63.13 62.42
PEPT 53.22 33.28 65.66 63.18 52.81 35.99 61.16 62.95 52.88 35.27 63.65 63.58
Dateset Es Biology Electronics
Method Metric MRR P@@@@1 P@@@@3 NDCG@@@@20 MRR P@@@@1 P@@@@3 NDCG@@@@20 MRR P@@@@1 P@@@@3 NDCG@@@@20
Doc2Vec 35.87 25.57 36.13 38.21 35.61 21.46 36.99 47.93 38.71 26.93 41.38 48.63
CNTN 38.97 27.20 40.78 46.93 33.58 20.17 33.69 36.91 49.52 27.40 65.38 61.46
NeRank 56.47 39.32 64.13 66.17 43.71 28.86 53.61 49.91 51.05 34.59 59.76 62.21
TCQR 47.16 30.59 59.41 60.53 39.62 24.06 4422 47.16 50.16 32.33 59.53 60.04
RMRN 55.16 37.61 63.04 64.53 46.62 31.53 55.62 55.78 56.22 40.38 67.59 67.29
UserEmb 47.03 30.32 57.35 48.69 34.97 22.63 36.75 46.72 41.75 29.54 45.41 51.70
EFHM 55.28 40.61 63.17 65.03 48.50 29.23 57.66 57.58 57.83 41.36 68.50 67.38
PMEF 57.35 42.69 65.20 66.61 47.19 32.16 56.93 56.88 58.72 42.38 69.50 68.29
ExpertBert 57.99 42.78 65.37 66.81 49.70 33.23 57.66 58.65 59.23 42.57 69.68 69.33
ExpertPLM 58.40 43.80 65.88 67.93 49.56 33.53 58.68 59.29 60.03 43.56 70.25 70.46
PEPT 59.66 44.76 66.97 68.38 51.06 34.53 60.00 62.13 61.23 44.70 71.33 71.98
Table 4 displays the overall results of our proposed model in comparison to other baselines. The experimental findings reveal several following observations:
  1. (1)
    Previous methods, such as Doc2Vec and CNTN, exhibit poor performance on almost all datasets. This attributes to the fact that these methods rely on the max or mean operator on experts’ histories to obtain expert features, which fails to consider the varying importance of different historical questions in modeling the interest of an expert.
  2. (2)
    In contrast, recent attention-based methods, such as RMRN and PMEF, have achieved better results. The reason is that these methods leverage attention mechanisms to selectively weigh the importance of different historical questions in modeling expert representations. It is worth noting that the PMEF outperforms most of the baselines. This is because PMEF utilizes a multi-view attentive paradigm, enabling it to learn multi-view representations of experts and questions.
  3. (3)
    Pre-training based methods (ExpertBert, ExpertPLM, PEPT) always outperform other methods. These methods employ pre-training technology to capture the general expert/question knowledge based on unsupervised data and transfer them to the expert finding. For example, ExpertPLM utilized a reputation-augmented MLM to capture the coarse-grained expert characteristics. And the model could capture the universal information from unlabeled data in a task-agnostic way, which can help learn better question-expert relatedness in expert finding.
  4. (4)
    As we can see, our proposed method PEPT surpasses two very recent SOTA work ExpertBert (Liu et al., 2022b) and ExpertPLM (Peng and Liu, 2022). The reason is that the ExpertBert could not capture expert abilities to answer different questions and the ExpertPLM could not capture more fine-grained question-level interest and expertise. And both of them are incapable of modeling personalized expert representations, which can reduce their ability to accurately model different experts.
In summary, the results validate the motivation proposed in Section 1 and reflect that PEPT with the characteristics in the Table 1 can achieve the best performance.
6.2. Ablation Study
Table 5. Characteristics of Different Variants.
Characteristics Variants Only Em w/o Expertise w/o Interest w/o Personalized PEPT
Expert-level Information $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
Vote Score $\times$ $\times$ $\surd$ $\surd$ $\surd$
ID Embedding $\times$ $\surd$ $\surd$ $\times$ $\surd$
MLM $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
Vote-oriented Task $\times$ $\times$ $\surd$ $\surd$ $\surd$
Question-level MLM $\times$ $\surd$ $\times$ $\surd$ $\surd$
To further explore how different modules of the pre-training model impact the model performance, we design diverse variants for experimenting (CodeReview and Gis datasets as examples), including:
  1. (1)
    Only Em, we adopt the original MLM task (Devlin et al., 2018) to further pre-train the bert-base-uncased model over the CQA expert-level corpus (i.e., one expert histories one input line);
  2. (2)
    w/o Expertise, we incorporate expert interest modeling and personalized information during the pre-training phase, and exclude the vote score information and vote-oriented task from our model;
  3. (3)
    w/o Interest, we introduce the expert expertise modeling and personalized information during pre-training, but remove the question-level masked language modeling task;
  4. (4)
    w/o Personalized, we incorporate the expert interest and expertise modeling but remove the personalized information during pre-training.
A general comparison of the four variants is provided in Table 5, and the experimental results are illustrated in Fig 3. Based on these results, we have made the following observations:
Figure 3. Ablation Study.
  1. (1)
    It is not surprising that Only Em obtains the worst performance. This is because the Only Em only uses expert-level word sequences (i.e., historical answered questions) and cannot effectively capture the question-level interest and ability required for finding suitable experts;
  2. (2)
    w/o Expertise and w/o Interest show varying degrees of decreased model performance. These results strongly reflect the importance of capturing expert interest and capability for finding suitable experts. Interestingly, the performance of w/o Interest is worse than that of w/o Expertise. We argue that interest may be more important than ability. In other words, experts in real life should first consider their interests rather than their abilities before answering questions;
  3. (3)
    Personalized information is crucial for modeling experts. As expected, removing the personalized information during pre-training leads to a decrease in model performance, which can be seen in w/o Personalized;
  4. (4)
    Our complete model PEPT achieves the highest performance among all the variants. The reason is that the PEPT is equipped with six characteristics in Table 5, which enable personalized pre-training of expert interest and expertise for one specific question.
Overall, the ablation studies provide evidence for our motivation and demonstrate the effectiveness of each component in our proposed model. Combining all key components in the PEPT framework leads to significant improvements in performance.
6.3. Pre-trained Expert Representation Analysis
Our pre-training paradigm aims to learn universal and initial expert representations from unsupervised data, which can be transferred to the downstream expert finding task. We conduct additional experiments aimed at exploring the influence of pre-trained expert representations.
6.3.1. Effect of Alleviating Data Sparsity
In real-world applications, expert finding often faces the challenge of data sparsity due to the need for amounts of supervised data. Our pre-training expert finding paradigm could have the potential to alleviate this problem. In this section, we investigate whether our pre-training paradigm can alleviate this problem. To simulate the scenario of data sparsity, we conduct experiments by varying the proportions of training datasets used during the fine-tuning process, namely $[100\%,80\%,60\%,40\%,20\%]$. Noted that the validation data and testing data used in these experiments are the same as those used in the main experiments. We select the PEPT, ExpertPLM, and PMEF for comparison.
Figure 4. Impacts of Fine-tuning Data Ratio. GAP = PEPT - PMEF.
Fig 4 shows the results on CodeReview and Electronics datasets. The performance of all models substantially drops when a smaller amount of training data is used. However, our pre-trained model, PEPT, consistently outperforms the best non pre-trained model, PMEF, in all cases. Furthermore, we notice a notable trend in the experimental results: as the amount of training data decreases, the performance gap between our proposed model, PEPT, and the baseline method, PMEF, becomes more pronounced. This observation suggests that the advantage of our pre-trained representations becomes even more pronounced when the training data is limited. This is because our pre-training paradigm enables the model to personalized leverage expert behaviors and vote scores for modeling experts’ interests and expertise. Besides, it could reduce the model’s dependence on training data.
Furthermore, we can find that although ExpertPLM also has the ability to alleviate data sparsity issues, it is still inferior to PEPT. This is because ExpertPLM can only pre-train coarse-grained expert representations, and it still needs sufficient supervised data for fine-tuning to capture specific representations of different experts. Different from that, PEPT can perform personalized expert representation learning during the pre-training stage and capture more fine-grained characteristics for expert finding. Overall, our experimental results suggest that PEPT can mitigate the impact of the data sparsity problem in expert finding.
Figure 5. Comparison with the Original BERT.
Figure 6. Impacts of Different Mask Ratio during Pre-training.
6.3.2. Comparison with the Original BERT
In our proposed method, the personalized expert pre-training paradigm is based on the original pre-trained BERT (i.e., bert-base-uncased). In this part, we further explore the effectiveness of this weight (derived by PEPT) by comparing it with the original BERT in the fine-tuning stage. Specifically, we replace our pre-trained weight with the BERT weight during the fine-tuning stage and compare the resulting performance.
The results are illustrated in Fig 5 (Gis and Biology as examples). Our experimental results indicate that the pre-trained weight obtained through our proposed pre-training paradigm is beneficial to the expert finding. As we noted previously, learning expert representations accurately is crucial for expert finding, as it involves capturing the interests and expertise of experts in answering various questions. However, the bert-base-uncased weight cannot capture these expert characteristics and can only find suitable experts based on text similarity matching. This observation highlights effectiveness of our pre-training paradigm in expert finding.
6.4. Further Analysis
In this section, we conduct experiments to evaluate the effect of key hyperparameters (including Masked Word Ratio, Masked Question Ratio, Learning Rate inf Fine-tuning) in our approach.
6.4.1. Masked Word Ratio
The masked word ratio is a critical hyperparameter in our proposed PEPT model during the expert representation pre-training. To investigate the sensitivity of parameters in the pre-training phase, we varied the ratio from $0.05$ to $0.25$ in increments of $0.05$.
The results are presented in Fig 6. All metric results initially improve as the mask ratio increases, reaching the optimal model performance at a certain point, and then gradually declining. If the word mask ratio is set too small, the pre-training process may not capture enough CQA language knowledge and result in reduced performance in the downstream task. Conversely, if the mask ratio is set too high, the [MASK] symbol may occur too frequently during pre-training, resulting in a mismatch between pre-training and fine-tuning, which could negatively impact performance. Hence, during the pre-training stage, a masked word ratio of $0.15$ is selected.
6.4.2. Masked Question Ratio
In our method, we design a question-level pre-training task for modeling expert interests. To explore the impacts of different masking question ratios on model performance, we varied the question masking ration in $[10\%,15\%,20\%,25\%,30\%]$ for experiments.
Figure 7. Impacts of Different Learning Rates during Fine-tuning.
Figure 8. Impacts of Different Historical Question Masking Ratios.
The results are presented in Fig 8. From the results, we find: (1) The ratio of masked questions can impact the model’s performance, with an initial increase followed by a decrease as the masking ratio increases. This is because if the question masking ratio is too small, the role of this pre-training task is weakened. And the ratio is too large could lead to a weakening of the model’s ability to accurately model experts; (2) Different datasets have different sensitivity to question mask ratios (e.g., the Biology dataset is $15\%$, the Electronic dataset is $20\%$). This is because the average number of expert answers in different datasets is also different (e.g., Biology dataset is $18.11$ while Electronic dataset is $33.14$). Hence, we need to set the question masking ratio carefully.
6.4.3. Learning Rate in Fine-tuning
Figure 9. Case Study. Visualization of matching scores between experts and questions. The arrows from experts to the same target question have the same color. Meanwhile, the thickness of the arrow represents the matching score (i.e., the thicker the arrow, the higher the matching score).
During fine-tuning, we observe an interesting phenomenon: the learning rate significantly impacts expert finding performance. As depicted in Fig 7, we observe a consistent pattern across different datasets: the model performance initially improves and then gradually declines. A learning rate that is too large can potentially destroy the expert representations captured by the model during pre-training, whereas a learning rate that is too small may not capture sufficient information for identifying experts on a specific dataset. Additionally, the inflection point of the curve differs across different datasets. For instance, the optimal learning rate for the English dataset is 5e-8, while that of the Biology dataset is 5e-6. We think that this might be due to the size of the dataset. For larger datasets (e.g., English), a higher learning rate could lead to drastic changes in the pre-trained representations, resulting in degraded expert finding performance. Conversely, smaller datasets (e.g., Biology) may require a higher learning rate to achieve optimal performance. In summary, we recommend setting the learning rate carefully.
6.5. Case Study
In this part, we conduct a case study to better understand what has been learned by the PEPT during pre-training. Specifically, we randomly select several target questions from the English testing datasets and several pre-trained candidate experts for prediction. We visualize the final matching score $S_{c}$ in Fig 9, where the thickness of the arrow represents the matching score. A thicker arrow indicates a higher matching score between questions and candidate experts.
As shown in the Figure, for the target question “Is data treated as … plural in formal context”, Expert A obtains the highest matching score and he/she is the ground truth expert. This is because Expert A has successfully answered relevant questions in the past and obtained a high vote score (e.g., ”Does staff take … plural verb” with a high vote score of +47). However, Expert D does not receive a high match score, even though he/she has answered a similar question (”One espresso, or … plural of espresso”). This is because his/her vote score is unsatisfactory (-1), indicating that he/she may not be suitable for answering this type question. It is not surprising that Expert C obtains the lowest matching score when matching any question. This is because Expert C obtained low vote scores in most historical questions, which might not be good for answering these questions.
These case studies demonstrate that our proposed paradigm can pre-train expert representations that include both interest and expertise, which can be transferred to the downstream task expert finding in CQA systems.
7. Conclusion
In this paper, our focus is on exploring the design of a personalized pre-training language model that is specifically customized for expert finding on CQA websites. Unlike the previous version ExpertPLM could only capture coarse-grained expert characteristics, our method is designed to precisely pre-train more personalized and fine-grained expert representations based on their histories. To achieve this goal, we propose a personalized pre-training framework that unifies expert historical answer questions, vote scores, and personalized information, then specially designs question-level masked language modeling and vote-oriented tasks. Based on these, the model is enabled to learn personalized expert features that can be transferred to the downstream task of expert finding, which could reduce the dependency on large amounts of labeled data. Our PEPT model demonstrates superiority expert finding performance on six real-world CQA datasets. Further ablation studies confirm the individual effectiveness of modeling interest, expertise, and personalization. Moreover, we conduct an evaluation to assess the performance of our model in scenarios with sparse training data. The results of this analysis highlight the model’s effectiveness in mitigating the issue of data sparsity.